OpenStudy (anonymous):

find the coordinates of the intersection of medians of the triangle whose vertices are (1,3), (-2,4) and (1,-2).

OpenStudy (shiraz14):

@kc_kennylau : Here's an interesting question to solve.

OpenStudy (kc_kennylau):

Shortcut: \(\Large\displaystyle\left(\frac{x_1+x_2+x_3}3,\frac{y_1+y_2+y_3}3\right)\) (I don't know if this is allowed)